# Orbital magnetoelectric coupling of three dimensional Chern insulators

### v2

@@ I.3 Layer Chern number
$(B_{3})=\sum_{s^{\prime}l^{\prime}}\left[D_{s^{\prime}l^{\prime},mm}(\textbf{k})-\frac{\sum_{\textbf{R}}\tilde{C}_{s^{\prime}l^{\prime},v}^{(0)\,*}(\textbf{R})\tilde{C}_{s^{\prime}l^{\prime},v}^{(0)}(\textbf{R})+\tilde{C}_{s^{\prime}l^{\prime},c}^{(0)\,*}(\textbf{R})\tilde{C}_{s^{\prime}l^{\prime},c}^{(0)}(\textbf{R})}{2}\right]D_{sl,n^{\prime}n}(\textbf{k})\frac{v_{\alpha,nm}(\textbf{k})v_{\beta,mn^{\prime}}(\textbf{k})}{\Delta_{\textbf{k},mn}\Delta_{\textbf{k},mn^{\prime}}}\,\tilde{\mathcal{E}}\,l^{\prime}.$ (46)
Then, the formula for $\alpha_{zz}^{(l)}$ is derived by singling out the electric field $\mathcal{E}$ and breaking the sum over layer index into parts for each layer, respectively, in Eq. (23).
@@
IV IV. Proof of layer index gauge invariance
We see from above that layer index appears explicitly in the expressions from $(B_{1})$ to $(B_{4})$. It seems that the results would depend on the gauge choice of layer index, which is a priori arbitrary. For example, if there are five layers, we can name them as $1,2,3,4,5$ or $-2,-1,0,1,2$ or in other similar ways. In this section, we will prove that the results of $\alpha_{zz}^{(l)}$ are independent of the gauge choice of layer index. Without loss of generality, this is equivalent to prove that the expression of $(B_{2})$, $(B_{3})$ and $(B_{4})$ is vanishing if setting $l^{\prime}=1$. Let us see what these terms become. For $(B_{2})$, we have
$(B_{2})$ $=(E_{m\textbf{k}}^{(0)}-\mu^{(0)})\Big{(}\sum_{\tilde{n}\neq n}\frac{\sum_{s^{\prime}l^{\prime}}D_{s^{\prime}l^{\prime},\tilde{n}n}(\textbf{k})D_{sl,n^{\prime}\tilde{n}}(\textbf{k})}{\Delta_{\textbf{k},n\tilde{n}}}+\sum_{\tilde{n}^{\prime}\neq n^{\prime}}\frac{\sum_{s^{\prime}l^{\prime}}D_{s^{\prime}l^{\prime},n^{\prime}\tilde{n}^{\prime}}(\textbf{k})D_{sl,\tilde{n}^{\prime}n}(\textbf{k})}{\Delta_{\textbf{k},n^{\prime}\tilde{n}^{\prime}}}\Big{)}$ (47)
@@
$-iA_{\alpha,nm}^{(1)\,*}$ $=\Big{[}\sum_{\tilde{n}\neq n}\frac{\sum_{s^{\prime}\,l^{\prime}}D_{s^{\prime}l^{\prime},n\tilde{n}}(\textbf{k})v_{\alpha,\tilde{n}m}(\textbf{k})}{\Delta_{\textbf{k},mn}\Delta_{\textbf{k},n\tilde{n}}}+\sum_{\tilde{n}\neq m}\frac{\sum_{s^{\prime}\,l^{\prime}}D_{s^{\prime}l^{\prime},\tilde{n}m}(\textbf{k})v_{\alpha,n\tilde{n}}(\textbf{k})}{\Delta_{\textbf{k},mn}\Delta_{\textbf{k},m\tilde{n}}}$ (48)
@@
$(B_{3})$ $=\big{[}\sum_{s^{\prime}l^{\prime}}D_{s^{\prime}l^{\prime},mm}(\textbf{k})-\frac{\sum_{s^{\prime}l^{\prime}\textbf{R}}\tilde{C}_{s^{\prime}l^{\prime},v}^{(0)\,*}(\textbf{R})\tilde{C}_{s^{\prime}l^{\prime},v}^{(0)}(\textbf{R})+\tilde{C}_{s^{\prime}l^{\prime},c}^{(0)\,*}(\textbf{R})\tilde{C}_{s^{\prime}l^{\prime},c}^{(0)}(\textbf{R})}{2}\big{]}$ (49)
@@
V V. Proof of the quantization rule in the interlayer decoupled case
In this section, we will prove rigorously the quantization rule Eq. (4) in the main text for the interlayer decoupled case. Then, the total Hamiltonian is obviously block-diagonal in the layer space. Each diagonal block is the Hamiltonian of every monolayer and they are all the same. If we diagonalize the full Hamiltonian, there are exactly $N_{\textrm{L}}$ degenerate energy valence bands and $N_{\textrm{L}}$ degenerate energy conduction bands if the slab consists of $N_{\textrm{L}}$ layers. Without loss of generality, we can associate valence band $n=1,\dots,N_{\textrm{L}}$ with layer index $l=1\dots N_{\textrm{L}}$, respectively. Similarly, we can also do the same for the conduction band $n=N_{\textrm{L}}+1,\dots,2N_{\textrm{L}}$ with layer index $l=1\dots N_{\textrm{L}}$, respectively. This leads to
$\sum_{s}D_{sl,n^{\prime}n}=\sum_{s}C_{sl,n^{\prime}}^{*}C_{sl,n}=\delta_{nn^{\prime}}\delta_{l,n/n-N_{\textrm{L}}}$ (50)
@@
$\sum_{s}(B_{2})$ $=E_{m\textbf{k}}^{(0)}\sum_{l^{\prime}}\Big{(}\sum_{\tilde{n}\neq n}\frac{\sum_{s^{\prime}}D_{s^{\prime}l^{\prime},\tilde{n}n}(\textbf{k})\sum_{s}D_{sl,n^{\prime}\tilde{n}}(\textbf{k})}{\Delta_{\textbf{k},n\tilde{n}}}+\sum_{\tilde{n}^{\prime}\neq n^{\prime}}\frac{\sum_{s^{\prime}}D_{s^{\prime}l^{\prime},n^{\prime}\tilde{n}^{\prime}}(\textbf{k})\sum_{s}D_{sl,\tilde{n}^{\prime}n}(\textbf{k})}{\Delta_{\textbf{k},n^{\prime}\tilde{n}^{\prime}}}\Big{)}$ (51)
@@
$\sum_{s}\sum_{n,n^{\prime}\in{\rm occ.}}\sum_{m\in{\rm emp}}(B_{4})$ $=\sum_{n\in{\rm occ.}}\sum_{m\in{\rm emp}}E_{m\textbf{k}}^{(0)}\,\delta_{l,n}\big{[}(-iA_{\alpha,nm}^{(1)})\frac{v_{\beta,mn}(\textbf{k})}{\Delta_{\textbf{k},mn}}+\frac{v_{\alpha,nm}(\textbf{k})}{\Delta_{\textbf{k},mn}(\textbf{k})}(iA_{\beta,nm}^{(1)*})\big{]}$ (52)
@@
If $m-N_{\textrm{L}}=n$, Eq. (52) equals 00 obviously. If $m-N_{\textrm{L}}\neq n$, since $v_{\alpha,nm}=\langle u_{n\textbf{k}}^{(0)}|(\partial_{k_{\alpha}}H_{\textbf{k}})|u_{m\textbf{k}}^{(0)}\rangle_{\Omega_{0}}$ is also block-diagonal, the expression must be zeros as well. To sum up, $(B_{4})$ also contributes 00.
@@
where we use a similar mapping as Eq. (50) for the system with OBC to derive the first equality. The last equality comes from the fact that $v_{\alpha,nm}=\langle u_{n\textbf{k}}^{(0)}|(\partial_{k_{\alpha}}H_{\textbf{k}})|u_{m\textbf{k}}^{(0)}\rangle_{\Omega_{0}}$ is also block-diagonal. Note that this expression is precisely the Chern number of monolayer. Therefore, if there are odd layers in system, the response is integer quantized; if $N_{\textrm{L}}$ is even, the respectively is then quantized to half-integer. Then, the quantization rule Eq. (4) in the main text is automatically proved for the interlayer decoupled case.
VI VI. Details on numerical evaluations and finite-size effect
In this section, we provide more details on the numerical calculations including the calculations of $\alpha_{zz}^{(l)}$ and layer Chern number.
@@ VI.1 Layer-resolved orbital magnetoelectric response of slab
In a slab system, we calculate $\alpha_{zz}^{(l)}$ using $\partial M_{z}^{(l)}/\partial\mathcal{E}$. We employ three approaches in our study. The first one is to calculate $M_{z}^{(l)}$ in the absence and in the presence of weak electric field using Eq. (23), then get $\alpha_{zz}^{(l)}$ by finite difference. In practice, we calculate $M_{z}^{(l)}$ for a series of weak electric field such that the response is still in the linear regime. Then, we fit the first four points to get the slope of the straight line and thus $\alpha_{zz}^{(l)}$. We use this for a slab with PBC in the $x,y$ plane.
Figure 2: Error in $\alpha_{zz}^{(l)}$ with respect to the quantized value using finite-difference method. (a) AA stacking, five layers; (b) AA stacking, six layers; (c) AB stacking, five layers; (a) AB stacking, six layers.
In Fig. 2, we show how the quantization is improved when we refine the k mesh. We see that the improvement is not monotonic but oscillating. The main source of error comes from the determination of chemical potential given by an auxillary system with OBC. Note that this is also the method we use to evaluate $\alpha_{zz}^{(l)}$ in the presence of substrate. More results for those shown in the main text are given in Table 1 where k mesh is $90\times 90$. We see that $\alpha_{zz}^{(l)}$ is not quantized in the presence of substrate and interlayer coupling because of inversion or mirror symmetry breaking.
Table 1: Layer-resolved response in units of [$e^{2}/h$] in systems with substrate.
@@
The second approach is to calculate $\alpha_{zz}^{(l)}$ directly applying perturbation formula given above. In Fig. 3, we show how the quantization is improved when we refine the k mesh. Compared to the first approach, one advantage of the perturbative approach is that the quantization gradually improves as the $k$ mesh becomes finer. We also note that another advantage of the perturbation approach is to enforce the mirror symmetry in the results such that $\alpha_{zz}^{(l)}$ is symmetric with respect to the central plane.
Figure 3: Error in $\alpha_{zz}^{(l)}$ with respect to the quantized value using perturbation theory. (a) AA stacking, five layers; (b) AA stacking, six layers; (c) AB stacking, five layers; (a) AB stacking, six layers.
The third approach is similar to the first approach except that we calculate $M_{z}^{(l)}$ using Eq. (4) instead of Eq. (23). This method is uniquely suitable for a slab with OBC in the $x,y$ plane. We use this method to test the robustness of the quantization rule against disorder.
@@ VI.2 Layer-resolved orbital magnetoelectric response of 3D bulk
In 3D bulk system, we calculate $\alpha_{zz}^{(l)}$ using $\partial P_{z}^{(l)}/\partial B$. However, the layer-resolved polarization $P_{z}^{(l)}$ is only well defined if the Wannier centers can be unambiguously associated with a given layer. We show in Fig. 4 that this is indeed the case. We see that the Wannier centers are always very close to the position of 2D layers.
Figure 4: Band structures of Wannier centers for all the cases we show in the main text: (a) AA stacking, five layers; (b) AA stacking, four layers; (c) AB stacking, four layers.
@@ VI.3 Layer Chern number
In the main text, we argue that the quantization rule of $\alpha_{zz}^{(l)}$ relies on quantized layer Chern number. In this subsection, we numerically check the quantization of layer Chern number using Eq. (25). In Fig. 5, we show the error of quantization for layer Chern number is gradually improved as k mesh becomes finer. Here, the interlayer coupling is $t_{3}=0.2$ and $t_{4}=0.15$ and the intralayer parameters are still the same as in the main text. Note that the error in layer Chern number is solely due to the finite mesh. The evaluation of layer Chern number using Eq. (25) does not reflect the topological character, therefore harder to converge compared to the usual way to calculate Chern number. For example, if we set interlayer coupling to be zero, the layer Chern number is exactly the 2D Chern number should be quantized as a topological invariant. However, for a fine mesh $100100\ times$ 100, the evaluation of Chern number using the plaquette method is quantized to precision $10^{-5}$, while the evaluation using Eq. (25) is only to precision $10^{-3}$.
Figure 5: Error in $C_{z}(l)$ with respect to the quantized value. (a) AA stacking, five layers; (b) AA stacking, six layers; (c) AB stacking, five layers; (a) AB stacking, six layers.
@@
Table 2: Layer Chern number $C_{z}(l)$ with substrate and those without substrate in the parenthesis.
@@
VII VII. Off-diagonal OME response of Chern insulators

### v3

@@ III Perturbation theory for layer-resolved orbital magnetoelectric response
$(B_{3})=\sum_{s^{\prime}l^{\prime}}\left[D_{s^{\prime}l^{\prime},mm}(\textbf{k})-\frac{\sum_{\textbf{R}}\tilde{C}_{s^{\prime}l^{\prime},v}^{(0)\,*}(\textbf{R})\tilde{C}_{s^{\prime}l^{\prime},v}^{(0)}(\textbf{R})+\tilde{C}_{s^{\prime}l^{\prime},c}^{(0)\,*}(\textbf{R})\tilde{C}_{s^{\prime}l^{\prime},c}^{(0)}(\textbf{R})}{2}\right]D_{sl,n^{\prime}n}(\textbf{k})\frac{v_{\alpha,nm}(\textbf{k})v_{\beta,mn^{\prime}}(\textbf{k})}{\Delta_{\textbf{k},mn}\Delta_{\textbf{k},mn^{\prime}}}\,\tilde{\mathcal{E}}\,l^{\prime}.$ (S46)
Then, the formula for $\alpha_{zz}^{(l)}$ is derived by singling out the electric field $\mathcal{E}$ and breaking the sum over layer index into parts for each layer, respectively, in Eq. (S23).
@@
IV Proof of layer index gauge invariance
We see from above that layer index appears explicitly in the expressions from $(B_{1})$ to $(B_{4})$. It seems that the results would depend on the “gauge choice” of layer index, which is a priori arbitrary. In other words, if there are five layers, we can name them as $1,2,3,4,5$ or $-2,-1,0,1,2$ or in other similar ways. In this section, we will prove that the results of $\alpha_{zz}^{(l)}$ are independent of the gauge choice of layer index. Without loss of generality, this is equivalent to prove that the expression of $(B_{2})$, $(B_{3})$ and $(B_{4})$ is vanishing if setting $l^{\prime}=1$. Let us see what these terms become. For $(B_{2})$, we have
$(B_{2})$ $=(E_{m\textbf{k}}^{(0)}-\mu^{(0)})\Big{(}\sum_{\tilde{n}\neq n}\frac{\sum_{s^{\prime}l^{\prime}}D_{s^{\prime}l^{\prime},\tilde{n}n}(\textbf{k})D_{sl,n^{\prime}\tilde{n}}(\textbf{k})}{\Delta_{\textbf{k},n\tilde{n}}}+\sum_{\tilde{n}^{\prime}\neq n^{\prime}}\frac{\sum_{s^{\prime}l^{\prime}}D_{s^{\prime}l^{\prime},n^{\prime}\tilde{n}^{\prime}}(\textbf{k})D_{sl,\tilde{n}^{\prime}n}(\textbf{k})}{\Delta_{\textbf{k},n^{\prime}\tilde{n}^{\prime}}}\Big{)}$ (S47)
@@ IV Proof of layer index gauge invariance
$-iA_{\alpha,nm}^{(1)\,*}$ $=\Big{[}\sum_{\tilde{n}\neq n}\frac{\sum_{s^{\prime}\,l^{\prime}}D_{s^{\prime}l^{\prime},n\tilde{n}}(\textbf{k})v_{\alpha,\tilde{n}m}(\textbf{k})}{\Delta_{\textbf{k},mn}\Delta_{\textbf{k},n\tilde{n}}}+\sum_{\tilde{n}\neq m}\frac{\sum_{s^{\prime}\,l^{\prime}}D_{s^{\prime}l^{\prime},\tilde{n}m}(\textbf{k})v_{\alpha,n\tilde{n}}(\textbf{k})}{\Delta_{\textbf{k},mn}\Delta_{\textbf{k},m\tilde{n}}}$ (S48)
@@
$(B_{3})$ $=\big{[}\sum_{s^{\prime}l^{\prime}}D_{s^{\prime}l^{\prime},mm}(\textbf{k})-\frac{\sum_{s^{\prime}l^{\prime}\textbf{R}}\tilde{C}_{s^{\prime}l^{\prime},v}^{(0)\,*}(\textbf{R})\tilde{C}_{s^{\prime}l^{\prime},v}^{(0)}(\textbf{R})+\tilde{C}_{s^{\prime}l^{\prime},c}^{(0)\,*}(\textbf{R})\tilde{C}_{s^{\prime}l^{\prime},c}^{(0)}(\textbf{R})}{2}\big{]}$ (S49)
@@
V Proof of the quantization rule in the interlayer decoupled case
In this section, we will prove rigorously the quantization rule Eq. (7) in the main text for the interlayer decoupled case, where the total Hamiltonian is obviously block-diagonal in the layer space. Each diagonal block is the Hamiltonian of every monolayer and they are all the same. If we diagonalize the full Hamiltonian, there are exactly $N_{\textrm{L}}$ degenerate energy valence bands and $N_{\textrm{L}}$ degenerate energy conduction bands if the slab consists of $N_{\textrm{L}}$ layers. Without loss of generality, we can associate valence band $n=1,\dots,N_{\textrm{L}}$ with layer index $l=1\dots N_{\textrm{L}}$, respectively. Similarly, we can also do the same for the conduction band $n=N_{\textrm{L}}+1,\dots,2N_{\textrm{L}}$ with layer index $l=1\dots N_{\textrm{L}}$, respectively. This leads to
$\sum_{s}D_{sl,n^{\prime}n}=\sum_{s}C_{sl,n^{\prime}}^{*}C_{sl,n}=\delta_{nn^{\prime}}\delta_{l,n/n-N_{\textrm{L}}}$ (S50)
@@ V Proof of the quantization rule in the interlayer decoupled case
$\sum_{s}(B_{2})$ $=E_{m\textbf{k}}^{(0)}\sum_{l^{\prime}}\Big{(}\sum_{\tilde{n}\neq n}\frac{\sum_{s^{\prime}}D_{s^{\prime}l^{\prime},\tilde{n}n}(\textbf{k})\sum_{s}D_{sl,n^{\prime}\tilde{n}}(\textbf{k})}{\Delta_{\textbf{k},n\tilde{n}}}+\sum_{\tilde{n}^{\prime}\neq n^{\prime}}\frac{\sum_{s^{\prime}}D_{s^{\prime}l^{\prime},n^{\prime}\tilde{n}^{\prime}}(\textbf{k})\sum_{s}D_{sl,\tilde{n}^{\prime}n}(\textbf{k})}{\Delta_{\textbf{k},n^{\prime}\tilde{n}^{\prime}}}\Big{)}$ (S51)
@@
$\sum_{s}\sum_{n,n^{\prime}\in{\rm occ.}}\sum_{m\in{\rm emp}}(B_{4})$ $=\sum_{n\in{\rm occ.}}\sum_{m\in{\rm emp}}E_{m\textbf{k}}^{(0)}\,\delta_{l,n}\big{[}(-iA_{\alpha,nm}^{(1)})\frac{v_{\beta,mn}(\textbf{k})}{\Delta_{\textbf{k},mn}}+\frac{v_{\alpha,nm}(\textbf{k})}{\Delta_{\textbf{k},mn}(\textbf{k})}(iA_{\beta,nm}^{(1)*})\big{]}$ (S52)
@@
If $m-N_{\textrm{L}}=n$, Eq. (S52) equals 00 obviously. If $m-N_{\textrm{L}}\neq n$, since $v_{\alpha,nm}=\langle u_{n\textbf{k}}^{(0)}|(\partial_{k_{\alpha}}H_{\textbf{k}})|u_{m\textbf{k}}^{(0)}\rangle_{\Omega_{0}}$ is also block-diagonal, the expression must be zeros as well. To sum up, $(B_{4})$ also contributes 00.
@@
where we use a similar mapping as Eq. (S50) for the system with OBC to derive the first equality. The last equality comes from the fact that $v_{\alpha,nm}=\langle u_{n\textbf{k}}^{(0)}|(\partial_{k_{\alpha}}H_{\textbf{k}})|u_{m\textbf{k}}^{(0)}\rangle_{\Omega_{0}}$ is also block-diagonal. Note that this expression is precisely the Chern number of monolayer. Therefore, if there are odd layers in system, the response is integer quantized; if $N_{\textrm{L}}$ is even, the respectively is then quantized to half-integer. Then, the quantization rule Eq. (7) in the main text is automatically proved for the interlayer decoupled case.
VI Details on numerical calculations and discussions on finite-size effect
In this section, we provide more details on the numerical calculations of $\alpha_{zz}^{(l)}$ and layer Chern number. We also discuss finite-size effect on the improvement of quantization.
@@ VI.1 Layer-resolved orbital magnetoelectric response of slab
In a slab system, we calculate $\alpha_{zz}^{(l)}$ using $\partial M_{z}^{(l)}/\partial\mathcal{E}$. We employ three approaches in our study. The used mathematical expressions are all given above.
The first approach is that we first calculate $M_{z}^{(l)}$ in the absence and in the presence of weak electric field using Eq. (2). Then, we get $\alpha_{zz}^{(l)}$ by finite difference. In practice, we calculate $M_{z}^{(l)}$ for a series of weak electric field such that the response is still in the linear regime. Then, we fit the results to get the slope of the straight line and thus $\alpha_{zz}^{(l)}$. This approach is uniquely suitable for a slab with OBC in the $x,y$ plane. We use this approach to test the robustness of the quantization rule against disorder.
The second approach is similar to the first one except using Eq. (S23), then get $\alpha_{zz}^{(l)}$ by finite difference. We use this for a slab with PBC in the $x,y$ plane. The obtained $\alpha_{zz}^{(l)}$ is quantized to integers if $N_{\rm L}$ odd or to half-integers if $N_{\rm L}$ even, as shown in Table S1. The distribution is symmetric to the central plane of the system and the total $\alpha_{zz}$ is indeed vanishing as dictated by mirror symmetry.
Table S1: Layer-resolved $\alpha_{zz}^{(l)}$ in units of $e^{2}/h$ obtained by numerical finite difference and those evaluated by perturbation theory given in the parenthesis. We consider $AA$,$AB$ stackings and $N_{\rm{L}}=5,6$.
Layer Index 1 2 3 4 5 6
Figure S2: Error in $\alpha_{zz}^{(l)}$ with respect to the quantized value using finite-difference method. (a) AA stacking, five layers; (b) AA stacking, six layers; (c) AB stacking, five layers; (a) AB stacking, six layers.
In Fig. S2, we show how the quantization is improved when we refine the k mesh. We see that the improvement is not monotonic but oscillating. The main source of error comes from the determination of chemical potential given by an auxillary system with OBC.
Note that this is also the method we use to evaluate $\alpha_{zz}^{(l)}$ in the presence of substrate. The obtained $\alpha_{zz}^{(l)}$ are given in Table S2 where k mesh is $90\times 90$. We see that $\alpha_{zz}^{(l)}$ is not quantized in the presence of substrate and interlayer coupling because of inversion or mirror symmetry breaking. However, the difference between neighboring $\alpha_{zz}^{(l)}$ is well quantized as shown in Table IV  of the main text.
Table S2: Layer-resolved response in units of $e^{2}/h$ in systems with substrate.
@@
The third approach is to calculate $\alpha_{zz}^{(l)}$ directly applying perturbation formula given above. In Fig. S3, we show how the quantization is improved when we refine the k mesh. Compared to the second approach, one advantage of the perturbative approach is that the quantization gradually improves as the $k$ mesh becomes finer. We also note that another advantage is to enforce the mirror symmetry in the results such that $\alpha_{zz}^{(l)}$ is symmetric with respect to the central plane.
Figure S3: Error in $\alpha_{zz}^{(l)}$ with respect to the quantized value using perturbation theory. (a) AA stacking, five layers; (b) AA stacking, six layers; (c) AB stacking, five layers; (a) AB stacking, six layers.
@@ VI.2 Layer-resolved orbital magnetoelectric response of 3D bulk
In 3D bulk system, we calculate $\alpha_{zz}^{(l)}$ using $\partial P_{z}^{(l)}/\partial B$. However, the layer-resolved polarization $P_{z}^{(l)}$ is only well defined if the Wannier centers can be unambiguously associated with a given layer. We show in Fig. S4 that this is indeed the case. We see that the Wannier centers are always very close to the position of 2D layers.
Figure S4: Band structures of Wannier centers for all the cases we show in the main text: (a) AA stacking, five layers; (b) AA stacking, four layers; (c) AB stacking, four layers.
@@ VI.3 Layer Chern number
In the main text, we argue that the quantization rule of $\alpha_{zz}^{(l)}$ relies on quantized layer Chern number. In this subsection, we numerically check the quantization of layer Chern number using Eq. (6). Here, the interlayer coupling is $t_{3}=0.2$ and $t_{4}=0.15$ and the intralayer parameters are still the same as in the main text. In Fig. S5, we show the error of quantization for layer Chern number is gradually improved as k mesh becomes finer. The error in layer Chern number is solely due to the finite mesh. The evaluation of layer Chern number using Eq. (6) does not reflect the topological character, therefore harder to converge compared to using the plaquette method to calculate Chern number. For example, if we set interlayer coupling to be zero, the layer Chern number is exactly the 2D Chern number should be quantized as a topological invariant. However, for a fine mesh $100\times 100$, the evaluation of Chern number using the plaquette method is quantized to precision $10^{-5}$, while the evaluation using Eq. (6) is only to precision $10^{-3}$.
Figure S5: Error in $C_{z}(l)$ with respect to the quantized value. (a) AA stacking, five layers; (b) AA stacking, six layers; (c) AB stacking, five layers; (a) AB stacking, six layers.
@@
Table S3: Layer Chern number $C_{z}(l)$ with substrate and those without substrate in the parenthesis.
@@
VII Off-diagonal OME response of Chern insulators